calculate the amount of energy required or realeased to cool down 12.5g of steam from 125.0 degree celsius to ice at -20 degree celsius


1
Expert's answer
2022-03-07T17:15:09-0500

Specific heat capacity of steam = Сsteam = 2.01 J g−1 °C−1

Specific heat capacity of water = Cwater = 4.186 J g−1 °C−1

Specific heat capacity of ice = Cice = 2.10 J g−1 °C−1

Latent heat of condensation of water = ∆Hcon = −2256 J/g

Latent heat of crystallization of water = ∆Hcry = −333.7 J/g


Solution:



This problem requires several steps:

Step 1: cool 12.5 g of steam from 125°C to 100°C

q1 = m × Сsteam × ∆T = (12.5 g) × (2.01 J g−1 °C−1) × (100°C − 125°C) = −628.125 J

q1 = −628.125 J


Step 2: convert 12.5 g of steam at 100°C to 12.5 g of water at 100°C

q2 = m × ∆Hcon = (12.5 g) × (−2256 J/g) = −28200 J

q2 = −28200 J


Step 3: cool 12.5 g of water from 100°C to 0°C

q3 = m × Cwater × ∆T = (12.5 g) × (4.186 J g−1 °C−1) × (0°C − 100°C) = −5232.5 J

q3 = −5232.5 J


Step 4: convert 12.5 g of water at 0°C to 12.5 g of ice at 0°C

q4 = m × ∆Hcry = (12.5 g) × (−333.7 J/g) = −4171.25 J

q4 = −4171.25 J


Step 5: cool 12.5 g of ice from 0°C to −20°C

q5 = m × Cice × ∆T = (12.5 g) × (2.10 J g−1 °C−1) × (−20°C − 0°C) = −525 J

q5 = −525 J


Add up all the heat required for all these changes is

q = q1 + q2 + q3 + q4 + q5

q = (−628.125 J) + (−28200 J) + (−5232.5 J) + (−4171.25 J) + (−525 J) = −38756.875 J = −38.76 kJ

q = −38.76 kJ

Thus, 38.76 kJ of energy realeased to cool down 12.5g of steam from 125°C to ice at −20°C


Answer: 38.76 kJ of energy realeased